Operations Report – May 5th

Crew 261 Operations Report 05-05-2023
SOL: 5
Name of person filing report: Julien Villa-Massone
Non-nominal systems:
· Remote observatory
· 1 Helmet, 1 headset
Notes on non-nominal systems: N/A
Spirit rover used: No
Hours: (before EVA) 220.5
Beginning charge: (Before EVA) 100
Ending charge: (On return from EVA, before recharging) N/A
Currently charging: Yes
Opportunity rover used: No
Hours: 123.0
Beginning charge: 100
Ending charge: N/A
Currently charging: Yes
Curiosity rover used: Yes
Hours: 233.7
Beginning charge: 100
Ending charge: 57
Currently charging: No (to save energy, as instructed by Mission Support)
Perseverance rover used: No
Hours: 263.1
Beginning charge: 100
Ending charge: N/A
Currently charging: Yes
General notes on rovers: Nominal operation today
Summary of Hab operations: Nominal day
WATER USE: rate of 28 gallons / sol today. Trend of 28 Gallons/sol since mission start. We are on target to finish the mission 11 gallons short (almost on target). This means, if we continue with this trend, the tank will be at 109 gallons when the mission ends.
Water (static tank): 328 gallons remaining
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: Nominal
Summary of suits and radios:
Headset #4 (previously erroneously reported as #1) is put aside for now, testing has not been possible yet due to time constraints.
Failed helmet has been set aside and labeled as such.
Summary of GreenHab operations: Nominal.
WATER USE: 6 gal at 06:30 + 6 gal at 19:00 + 1 gal science = 13 gals total
Heater: Off
Supplemental light: Off
Harvest: 110g red tomatoes, 340g cucumber, 600g orange tomatoes, 20g mint.
Summary of ScienceDome operations: Nominal.
Dual split: Off
POWER report: Nominal operation, except for a breaker disconnected between a PV string and an MPPT. It is the third occasion this has been observed since mission start – it has been the same breaker every time. Reset as per usual.
Summary of RAM operations: Crew Robotics Engineer made progress by attaching the sails of the robot and performing various other adjustments, due for testing soon.
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nominal.
Questions, concerns and requests to Mission Support: Thank you for supporting us today.

Copyright © The Mars Society. All rights reserved. | Main Site